Al can paint a certain house in 10 days and Ben can paint the house in 12 days. How long will it take to paint the house if both men work?
for this one I got
1/12+1/10 common dom 120 and 22/120 than simplify to 11/60 so for both men it would take 11 days

Answer by scott8148(6628)

you're on the right track, but how can it take longer than the fastest man?

if it takes x days, then Al's portion is x/10 and Ben's portion is x/12

together they do the whole job __ x/10 + x/12 = 1

multiplying by 60 __ 6x + 5x = 60 __ 11x = 60 __ x = 60/11

what you found was one day's portion of the job